120SnD_{4}: A SYMMETRIC TOP?

\underline{M. HALONEN}, and L. HALONEN, Laboratory of Physical Chemistry, P. O. Box 55 (A. I. Virtasen aukio 1), FIN-00014 University of Helsinki, Finland; H. BÜRGER and W. JERZEMBECK, FB9-Anorganische Chemie, Universität-GH Wuppertal, D-42097 Wuppertal, Germany.